# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v2

@@ 4.5 Extended evaluation of unsupervised multimodal change detection methods
Limitations in traditional UMCD datasets: The traditional UMCD datasets consist of only a handful of image pairs, often depicting simple land-cover changes such as urban expansion or deforestation, with relatively low-resolution imagery with limited geographic diversity, as shown in Figure 12. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, BRIGHT provides thousands of high-resolution multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
@@
Interference from non-disaster changes: Unlike previous UMCD benchmarks where general land-cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, BRIGHT requires methods to differentiate true building structural damage from irrelevant land-cover variations, posing a unique challenge to current approaches.
@@
Problematic evaluation protocol: Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real disaster scenario, there are no labeled samples available for hyperparameter tuning, making the current practice unrealistic. BRIGHT exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results highlight the challenges of applying existing UMCD models to real-world disaster scenarios. The BRIGHT dataset reveals significant limitations in current methodologies and also presents new opportunities for future UMCD research.
@@ 4.6 Extended evaluation of unsupervised multimodal image matching methods
Table 12 presents the quantitative results of several representative UMIM methods on the Noto-Earthquake-2024 scene. Our experimental results reveal that traditional feature-based methods generally fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable. In contrast, area-based methods, which rely on region-level similarity rather than point-level correspondence, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that the BRIGHT benchmark and its challenging real-world settings will inspire the development of new UMIM methods that are more robust to large scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
@@ 5.1 Limitation of BRIGHT
The performance of ML/DL models heavily depends on the quality of the training data. The BRIGHT dataset is designed to enable ML/DL models trained on it to generate building damage proxy maps of practical significance for future disaster events. It is anticipated that BRIGHT will serve as a benchmark for numerous subsequent studies and practical disaster relief applications.
However, the BRIGHT dataset also has some limitations. Firstly, our dataset involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 4. Secondly, the building polygons in BRIGHT were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on BRIGHT. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise. Finally, although BRIGHT is globally distributed and richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 v.s. Hawaii-Wildfire-2023 in Table 2. Although we used an additional event-wise evaluation method, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. It may affect the generalizability of the trained model in real disaster scenarios.
Overall, despite the limitations, it is the first time such a multimodal VHR dataset has been constructed that caters to a transferable approach to obtain building damage proxy maps at different geographical locations.
@@ 5.2 Significance of BRIGHT
Latency in responding to disaster events often arises from delays in interpretation methods and the acquisition of EO data (Ye et al., 2024). Traditionally, building damage proxy maps are generated through expert visual interpretation, a time-consuming process. Although ML- and DL-based methods offer automated and efficient approaches to building damage proxy mapping, they are often limited by the available training data. Large open-source optical datasets, though valuable, restrict models to optical imagery alone. The acquisition of clean optical EO data also requires sunlight and a cloud-free environment, which can result in delays of several days for certain disaster events such as typhoons and wildfires. As the first globally distributed multimodal dataset, BRIGHT encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on BRIGHT to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, BRIGHT offers several distinct characteristics: multimodal data, VHR imagery with resolutions better than 1 meter per pixel, coverage of six types of natural disasters, as well as human-made disasters, rich geographic diversity, and open access to the community. Due to these features, BRIGHT is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications. BRIGHT facilitates research on building damage assessment. Looking forward, BRIGHT can be expanded to include additional data modalities, such as height data, which would provide valuable topographic information to enhance disaster response strategies. Incorporating more bands in optical and SAR data could also broaden the applicability of the dataset.
In addition to its primary application in building damage assessment, the BRIGHT dataset holds significant potential for other domains, particularly in multimodal EO data-related research. Firstly, BRIGHT can be directly used as a building detection benchmark dataset with a rich geographic diversity and building patterns. Beyond that, researchers can also leverage BRIGHT to generate new datasets tailored to specific needs to explore new techniques including, but not limited to, image registration, land cover change detection, land cover mapping, height estimation, EO-driven visual question answering (VQA), improving the performance and reliability of models in complex environments. Moreover, BRIGHT is a potential resource for the development of remote sensing foundation models (Wang et al., 2023; Hong et al., 2024). These emerging models, designed to generalize across various tasks and datasets, benefit immensely from the rich multimodal data and extensive geographic diversity offered by BRIGHT. By incorporating both optical and SAR data, BRIGHT enables the pre-training of more robust foundation models capable of handling the complexities of real-world remote sensing applications. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024). We envision that BRIGHT, true to its name, will bring even a glimmer of bright to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
@@
A prompt and comprehensive understanding of building damage following a disaster is critical for effective disaster response and relief operations. Currently, large-scale optical datasets support the development of automated ML/DL methods, but their use is limited by weather conditions and daylight availability. In contrast, all-weather disaster response can be facilitated by leveraging multimodal EO data. In this paper, we introduced BRIGHT, the first globally distributed, multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. BRIGHT includes pre-event optical imagery and post-event SAR imagery with very high spatial resolutions (better than 1 meter). Alongside the dataset, we evaluated several state-of-the-art models on BRIGHT. The experimental results demonstrate the effectiveness of BRIGHT in mapping building damage across unseen events. These findings provide both performance baselines and valuable insights into model design for future research. BRIGHT is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve BRIGHT’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.
@@ Appendix A Feature visualization
To visualize the spatial focus of models during building damage mapping, we apply Grad-CAM (Selvaraju et al., 2017), a gradient-based technique for generating class-specific activation maps. We perform Grad-CAM analysis on the ChangeOS model. Specifically, we compute the activation responses for the “Damaged” and “Destroyed” categories from the following components: the third or fourth ResNet block of the two branches of encoder, capturing mid- to high-level modality-specific features; the fusion module of the damage classification branch in the decoder, reflecting early-stage fusion and decision-making. These activations help reveal which spatial regions the model relies on when assessing different damage levels and how it utilizes information from optical and SAR modalities differently depending on the disaster context.
@@ Appendix B Manual registration and estimating registration errors
Despite multiple rounds of meticulous registration and cross-validation of BRIGHT by several EO experts, registration errors between optical and SAR images could not be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we adopt a proxy method for registration error estimation:
@@
Feature Point Selection: Select many feature points (e.g., 3000 points) from both the optical and SAR images using a keypoint detection algorithm.
@@
Figure of selected feature points (can be used as proxy ground truth for evaluating descriptors)
@@ Appendix C Details of benchmark DL models
For DeepLabV3+, we applied ResNet-50 initialized with ImageNet pretraining weights (He et al., 2016) as the encoder. We then modified the input channel of the encoder so that the network can predict damage proxy maps directly from the optical and SAR stacked images.
@@ Appendix D Details of post-processing methods
To further refine the raw predictions from deep learning models and enhance the quality of damage proxy maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 4.3. This section provides an overview of these methods and details the specific implementation used in our study.
@@
Object-based post-processing aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
@@ Appendix E Details of unsupervised domain adaptation methods
To evaluate the effectiveness of unsupervised domain adaptation (UDA) techniques in the cross-event zero-shot setting, we selected several representative methods from the computer vision literature originally designed for semantic segmentation tasks. The following methods are implemented:
@@
In our experimental setup, each test disaster event is treated as the target domain, and only its unlabeled optical-SAR image pairs are used for adaptation. Importantly, no labeled target domain data is used for training or model selection, in line with our zero-shot assumption. This setup is fundamentally different from conventional UDA protocols in the computer vision community, where a target domain validation set is typically available to select the best model checkpoint, and the target test set is held out purely for evaluation. In contrast, our setting is designed to reflect real-world disaster response scenarios, where it is realistic to obtain unlabeled imagery, but no ground-truth annotations are available at adaptation time. Therefore, model selection must rely solely on source-domain feedback, which introduces additional challenges and more closely aligns with the operational constraints of emergency response.
@@
For all methods, the same data preprocessing pipeline and training schedule (e.g., learning rate, batch size) are used unless otherwise specified in the original method, like those supervised models. The final evaluation is conducted on the target domain’s test set, and performance is reported in terms of mIoU.
@@ Appendix F Details of semi-supervised learning methods
Appendix G Introduction of UMCD methods and a more practical evaluation protocol
We further evaluate several recent UMCD methods on the BRIGHT dataset. These methods are designed to detect binary changes from paired multimodal imagery without relying on manual labels.

### v3

@@ 4.5 Extended evaluation of unsupervised multimodal change detection methods
Limitations in traditional UMCD datasets. The traditional UMCD datasets consist of only a handful of image pairs, often depicting simple land-cover changes such as urban expansion or deforestation, with relatively low-resolution imagery with limited geographic diversity, as shown in Figure 12. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, Bright provides thousands of VHR multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
@@
Interference from non-disaster changes. Unlike previous UMCD benchmarks where general land-cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, Bright requires methods to differentiate true building structural damage from irrelevant land-cover variations, posing a unique challenge to current approaches.
@@
Problematic evaluation protocol. Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real disaster scenario, there are no labeled samples available for hyperparameter tuning, making the current practice unrealistic. Bright exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results highlight the challenges of applying existing UMCD models to real-world disaster scenarios. Bright reveals significant limitations in current methodologies and also presents new opportunities for future UMCD research.
@@ 4.6 Extended evaluation of unsupervised multimodal image matching methods
Table 12 presents the quantitative results of four UMIM methods on the Noto-Earthquake-2024 scene. These methods fall into two main categories: feature-based methods (LNIFT (Li et al., 2022) and SRIF (Li et al., 2023a)), which rely on sparse keypoint detection and matching, and area-based methods (FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019)), which operate on local regions to estimate correspondences. A detailed description of each method is provided in Appendix H. Our experimental results reveal that traditional feature-based methods generally fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable. In contrast, area-based methods, which rely on the similarity in a local region rather than point correspondence in the global level, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that the Bright benchmark and its challenging real-world settings will inspire the development of new UMIM methods that are more robust to large scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
@@ 5.1 Limitation of Bright
The performance of ML/DL models heavily depends on the quality of the training data. The Bright dataset is designed to enable ML/DL models trained on it to generate building damage maps of practical significance for future disaster events. It is anticipated that Bright will serve as a benchmark for numerous subsequent studies and practical disaster relief applications. However, the Bright dataset also has some limitations.
1.
Registration error. Bright involves both optical and SAR imagery covering the same locations. SAR images, in particular, can be distorted and stretched in certain areas. Despite thorough preprocessing, including manual alignment and cross-checking by multiple experts in EO data processing, minor alignment errors may still persist, as suggested by Table 4.
2.
Label quality. The building polygons in Bright were manually annotated by expert annotators. Although manual labeling generally ensures high accuracy, minor errors in polygon boundaries are inevitable due to the complexity of building shapes and the variability in image resolution. These inaccuracies may slightly affect the performance of the models trained on Bright. Then, the extent of damage to buildings is assessed by experts through visual interpretation of optical EO data. This process is susceptible to occasional misjudgments, contributing to label noise.
3.
Sample and regional imbalance. Although Bright is richly geographically diverse, it has the problem of regional imbalance in the number of labels. Some of its events have more tiles and building numbers and thus are more dominant in training and evaluation, e.g., Turkey-Earthquake-2023 (1,114 tiles) v.s. Hawaii-Wildfire-2023 (65 tiles) in Table 2. Although we used an additional event-level evaluation method, the effect of events, which account for a large percentage of the sample, on the model during the training phase is still not negligible. It may affect the generalizability of the trained model in real disaster scenarios. Apart from that, all disaster events are located near the equator or in the northern hemisphere, with no events from the southern hemisphere represented. This spatial bias could potentially limit the applicability of trained models to regions with different building styles, land cover patterns, or SAR imaging characteristics prevalent in the southern part of the globe.
Overall, despite the limitations, it is the first time such a multimodal VHR dataset has been constructed that caters to multimodal EO research under large-scale and diverse disaster context.
@@ 5.2 Significance of Bright
Timely disaster response is often hindered by delays in both EO data acquisition and damage interpretation workflows (Ye et al., 2024). Traditional expert-driven building damage mapping is time-consuming and not scalable. While ML and DL methods offer automated alternatives, their effectiveness remains limited by the scope and quality of available training data. Most existing open-source datasets are optical-only, restricting the models’ operational applicability in adverse weather and low-light conditions. As the first globally distributed multimodal dataset, Bright encompasses both pre-event optical imagery and post-event SAR imagery. This unique combination overcomes the limitations of optical EO data by enabling models trained on Bright to monitor disaster-stricken areas regardless of weather conditions or daylight. Compared to existing building damage datasets, Bright offers several distinct characteristics: multimodal data, VHR imagery with sub-meter spatial resolution, coverage of five types of natural disasters, as well as two kinds of human-made disasters, rich geographic diversity, and open access to the community. Due to these features, Bright is anticipated to serve as a benchmark for many subsequent studies and practical disaster relief applications.
Beyond building damage assessment, Bright can also be used to support several research directions within the EO and vision community. In this work, we have applied Bright to evaluate the performance of several UDA and SSL methods for cross-event transfer. Additionally, we have demonstrated its applicability to UMCD and UMIM, showcasing its versatility as a benchmark for multiple EO challenges under real-world constraints. Furthermore, the dataset provides a strong foundation for broader multimodal EO research. Its high-quality annotations and geographic diversity make it directly applicable to tasks such as building footprint extraction, land cover mapping, height estimation, and even EO-based visual question answering (VQA). Researchers can also repurpose or extend Bright to create task-specific benchmarks, enabling flexible experimentation across tasks. Bright is also well-positioned to support the development of remote sensing foundation models, large-scale, pre-trained models designed to generalize across sensors, tasks, and regions (Wang et al., 2023; Hong et al., 2024). Its rich combination of modalities, spatial detail, and contextual diversity provides the kind of data diversity required to build such general-purpose, transferable models. This contribution is significant as the field moves toward creating versatile, scalable AI models that can be applied across different types of EO data and disaster scenarios (Li et al., 2024)
Looking ahead, Bright can be further enhanced by incorporating additional modalities, such as elevation data or multi-band SAR/optical inputs, to support terrain-aware analysis and improve semantic reasoning. Future versions may also help fill current geographic gaps, including southern hemisphere regions, ensuring more globally representative coverage. Ultimately, We envision that Bright, true to its name, will bring even a glimmer of brightness to people in disaster-stricken areas by enabling more prompt and effective disaster response and relief.
@@
In this paper, we introduced Bright, the first globally distributed, multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. Bright includes pre-event optical imagery and post-event SAR imagery with sub-meter spatial resolution. Beyond introducing the dataset, we conducted a comprehensive series of experiments to validate its utility. We benchmarked several state-of-the-art supervised learning models under a standard machine learning split, and further extended the evaluation to a cross-event transfer setting, simulating real-world scenarios where no/limited target annotations are available. Additionally, we assessed the performance of unsupervised domain adaptation and semi-supervised learning methods, as well as unsupervised multimodal change detection and image matching techniques. These findings provide both performance baselines and valuable insights into model design for future research. Bright is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve Bright’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.
@@ Appendix A Feature visualization
To visualize the spatial focus of models during building damage mapping, we apply Grad-CAM technique999https://github.com/jacobgil/pytorch-grad-cam, a gradient-based technique for generating class-specific activation maps (Selvaraju et al., 2017). We perform Grad-CAM analysis on the ChangeOS model. Specifically, we compute the activation responses for the “Damaged” and “Destroyed” categories from the following components: the third or fourth ResNet block of the two branches of encoder, capturing mid- to high-level modality-specific features; the fusion module of the damage classification branch in the decoder, reflecting early-stage fusion and decision-making. These activations help reveal which spatial regions the model relies on when assessing different damage levels and how it utilizes information from optical and SAR modalities differently, depending on the disaster context.
@@ Appendix B Manual registration and estimating registration errors
The manually selected control points by EO experts on some disaster scenes are shown in Figure 11. Despite multiple rounds of meticulous registration and cross-validation by several EO experts, registration errors between optical and SAR images cannot be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we propose a proxy method for registration error estimation:
@@
Feature point selection: Select many feature points (e.g., 3000 points) from both the optical and SAR tiles using a keypoint detection algorithm.
@@ Appendix C Details of benchmark DL models
For DeepLabV3+, we applied ResNet-50 initialized with ImageNet pretraining weights (He et al., 2016) as the encoder. We then modified the input channel of the encoder so that the network can predict building damage maps directly from the optical and SAR stacked images.
@@ Appendix D Details of post-processing methods
To further refine the raw predictions from deep learning models and enhance the quality of building damage maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 4.3. This section provides an overview of these methods and details the specific implementation used in our study.
@@
Object-based major voting aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
@@ Appendix E Details of unsupervised domain adaptation methods
To evaluate the effectiveness of UDA techniques in the cross-event zero-shot setting, we selected several representative methods from the computer vision literature originally designed for semantic segmentation tasks. The following methods are implemented:
@@
In our experimental setup, each test disaster event is treated as the target domain, and only its unlabeled optical-SAR image pairs are used for adaptation. Importantly, no labeled target domain data is used for training or model selection, in line with our zero-shot assumption. This setup is fundamentally different from conventional UDA protocols in the computer vision community, where a target domain validation set is typically available to select the best model checkpoint, and the target test set is held out purely for evaluation. In contrast, our setting is designed to reflect real-world disaster response scenarios, where it is realistic to obtain unlabeled imagery, but no ground-truth annotations are available at adaptation time. Therefore, model selection must rely solely on source domain feedback, which introduces additional challenges and more closely aligns with the operational constraints of emergency response.
@@
For all methods, the same data preprocessing pipeline and training schedule (e.g., learning rate, batch size) are used unless otherwise specified in the original method, like those supervised models. The final evaluation is conducted on the target domain, and performance is reported in terms of mIoU.
@@ Appendix F Details of semi-supervised learning methods
Appendix G Details of unsupervised multimodal change detection methods and a more practical evaluation protocol
We further evaluate several recent UMCD methods on Bright. These methods are designed to detect binary changes from paired multimodal imagery without relying on manual labels.
@@ Appendix G Details of unsupervised multimodal change detection methods and a more practical evaluation protocol
Appendix H Details of unsupervised multimodal image matching methods
In our extended evaluation of UMIM methods, we categorized the selected algorithms into two groups based on their matching strategies: feature-based and area-based methods.
Feature-based methods, such as LNIFT (Li et al., 2022) and SRIF (Li et al., 2023a), aim to detect sparse keypoints independently in both optical and SAR images and compute modality-invariant feature descriptors. The matching is then performed globally by comparing descriptor similarity between the two modalities. These methods are implemented using publicly available code101010https://github.com/LJY-RS/SRIF.
In contrast, area-based methods, including FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019), begin by identifying keypoints only in the optical image. These points are then matched by sliding a template window across a local search region in the SAR image, relying on local patch similarity. In our experiments, we used a template size of 120120 pixels, and a search window of 200200 pixels for both methods. This category is also implemented using open-source code111111https://github.com/yeyuanxin110/CFOG.
We performed registration on the entire original EO image pair from the Noto-Earthquake-2024 event. Because of the large spatial size of the input images, images were downsampled to half their original resolution to improve computational efficiency and maintain feasibility for all methods.
For quantitative evaluation, we adopted a control-point-based proxy metric. A set of manually selected control points, validated by EO experts shown in Figure 11, is used to represent stable, cross-modal, and clearly identifiable features (e.g., building corners, road intersections). The average spatial offset is computed using their coordinates in a common projected coordinate system (e.g., UTM), both before and after registration. Before registration, the offset can be calculated as:
where and are the projected coordinates of the -th control point in the optical and SAR images, respectively.
After registration, SAR points are transformed using the estimated mapping , and the post-registration error is computed as:
where are transformed SAR coordinates.
This approach provides an alternative approximation of alignment accuracy in the absence of ground truth. Although this metric is not a substitute for full correspondence maps, it offers a practical and interpretable measure of registration performance, and enables us to assess how well unsupervised methods can approach human-level matching under complex, real-world disaster conditions.